A population consists of the following measurement 3,8,14,17,and 58 consider all the sample of size 3 taken from this population without replacement


(1)List the sample of the mean and media respectively
(2) calculate the mean of the sample distribution of the media and the means
1
Expert's answer
2021-04-14T09:03:59-0400

Given sample- {3,8,14,17,58}

n=3


(1) The List of samples are provided in the table below-




(2) Mean of sample mean "=\dfrac{\sum \text{ Sample mean}}{\text{size of sample}}"


"=\dfrac{186.97}{9}=20.77"